# Smartify: A Multi-Agent Framework for Automated Vulnerability Detection and Repair in Solidity and Move Smart Contracts

### v1

@@ Abstract
The rapid growth of the blockchain ecosystem and the increasing value locked in smart contracts necessitate robust security measures. While languages like Solidity and Move aim to improve smart contract security, vulnerabilities persist. This paper presents Smartify, a novel multi-agent framework leveraging Large Language Models (LLMs) to automatically detect and repair vulnerabilities in Solidity and Move smart contracts. Unlike traditional methods that rely solely on vast pre-training datasets, Smartify employs a team of specialized agents working on different specially fine tuned LLMs to analyze code based on underlying programming concepts and language-specific security principles. We evaluated Smartify on a dataset for Solidity and a curated dataset for Move, demonstrating its effectiveness in fixing a wide range of vulnerabilities. Our results show that Smartify (Gemma2+codegemma) achieves state-of-the-art performance, surpassing existing LLMs and even enhancing the capabilities of general-purpose models, such as Llama 3.1. Notably, Smartify can incorporate language-specific knowledge, such as the nuances of Move, without requiring massive language-specific pre-training datasets. This work offers a detailed analysis of various LLMs’ performance on smart contract repair, highlighting the strengths of our multi-agent approach and providing a blueprint for developing more secure and reliable decentralized applications in the growing blockchain landscape. We also provide a detailed recipe for extending this to other similar use cases.
Keywords S
mart Contracts, Vulnerability Detection, Code Repair, Large Language Models, Blockchain Security, Move, Solidity
@@ 1 Introduction
Smart contracts, self-executing agreements with terms directly written into code, have emerged as a cornerstone of blockchain technology [1, 2]. Their ability to automate transactions and eliminate intermediaries has led to widespread adoption in various sectors, including finance, supply chain management, and healthcare [3, 4, 5]. However, the increasing complexity of smart contracts has given rise to a growing concern: security vulnerabilities[6]. These vulnerabilities, often stemming from coding errors or design flaws, can be exploited by malicious actors, leading to significant financial losses and damage to the reputation of blockchain projects.
@@
The financial implications of smart contract vulnerabilities are substantial. Reports indicate that cumulative losses from attacks against Ethereum smart contracts alone have exceeded USD 3.1 billion as of 2023 [7]. In the DeFi space, an estimated \$9.04 billion has been stolen due to vulnerabilities [8]. Notable incidents like the DAO hack of 2016, resulting in a \$55 million loss [9], and the Poly Network hack in 2021, where over \$600 million was stolen [10], underscore the critical need for robust security measures.
@@
Traditional security auditing methods, while essential, often face limitations in terms of accuracy and scalability. This has spurred the exploration of automated techniques for vulnerability detection [11, 12]and repair, with Large Language Models (LLMs) emerging as a promising solution [13]. LLMs, trained on vast datasets of code, can learn to understand and generate code that adheres to specific programming paradigms and best practices. However, most of the the tools available for smart contracts are very language-specific, mostly relying on Solidity as the language of choice, as well as often sometimes requiring compiled bytecode for scanning for other languages[14].
@@
Apart from Solidity [15], Move [16] has gained significant traction lately due to its strong focus on security. Its cutting-edge features, including a custom data type for secure operations and robust access controls via Move modules, and unique memory safety features [17] have been particularly noteworthy. Moreover, the Move Prover, a native security framework, provides an additional layer of protection [18]. Notably, several prominent blockchain platforms, such as Starcoin [19], Aptos [20], and Sui [21], have already adopted Move.
@@
However, despite its promising architecture, the real-world security performance of Move modules remains largely untested. Unlike Solidity-based smart contracts, which have been extensively studied through empirical research and surveys, there is a scarcity of research focused specifically on Move modules. Although some methodologies have been proposed for identifying defects in Move modules or conducting formal verification [22, 23], and empirical analysis[14], a significant knowledge gap persists. Specifically, large-scale investigations into the frequency of defects in real-world Move modules and identifying potential vulnerabilities and repairing them are lacking, highlighting the need for further research in this area.
This paper proposes a novel framework for detecting and repairing vulnerabilities in smart contracts, focusing on the Solidity and Move languages from a programming language perspective. Our hypothesis relies on understanding the code and preventing known bad practices and unsafe code from being written before even compilation to prevent vulnerability. Our approach leverages the power of a multi-LLM agent system, combining the strengths of explanation and repair models. Our framework, Smartify, leverages a multi-agent LLM framework to understand, critique, and repair code based on previously learned vulnerabilities as well as propose patches to repair them. By integrating an LLM specialized in code explanation with another focused on code repair, we aim to improve the accuracy and efficiency of the vulnerability remediation process.
@@
We try to answer the following research questions in this paper, related to software engineering using AI agents and in the landscape of complex smart contract reasoning.
@@
  • RQ4: Does the proposed post-training framework be generalized to a larger set of pre-trained LLMs?
The key contributions of this paper are as follows:
  1. 1.
    We introduce Smartify, a multi-agent LLM code detection and repair framework that can analyze and repair codes based on coding concepts instead of just using the vast amount of codes for pre-training.
  2. 2.
    We propose a method that can encompass programming language-specific paradigms for smart contracts, both for established language like Solidity and low resource language like Move, without the need for significantly large pertaining dataset.
  3. 3.
    We give a detailed recipe for how this can be scaled for other languages and give a comprehensive evaluation of Smartify’s efficacy for other pre trained LLMs.
  4. 4.
    We introduce, implement, and evaluate our framework on generalized pre trained LLMs to show the efficacy of our framework. We evaluate the performance of our framework and various LLMs on a diverse set of vulnerabilities in Solidity and Move smart contracts.
  5. 5.
    We provide a detailed analysis of the results, identifying the strengths and weaknesses of different approaches and highlighting the challenges in automated code repair.
@@ 2 Related Work
This section reviews related work in smart contract vulnerabilities, security auditing tools, traditional code repair techniques, and the emerging use of Large Language Models (LLMs) for code repair, particularly in the context of Solidity and Move.
2.1 Smart Contract Vulnerabilities
Smart contracts, while offering automation and trustless execution, are prone to security vulnerabilities due to their complex code, immutable nature, and the decentralized environment they operate in [24, 25, 26]. Exploiting these vulnerabilities can lead to severe financial losses, service disruptions, and loss of trust in decentralized applications [27]. Common vulnerabilities include:
Reentrancy: This occurs when a malicious contract calls back into the original contract before the first function invocation completes[28, 29]. This can disrupt control flow, allowing attackers to repeatedly execute a vulnerable function, potentially draining funds or manipulating the contract’s state [29, 30].
Integer Overflow/Underflow: These vulnerabilities arise when arithmetic operations result in values exceeding the maximum or falling below the minimum representable value for the integer type. Before Solidity 0.8.0, these errors wrapped around silently, leading to unexpected behavior.
Access Control Issues: Insufficient or improperly implemented access control can allow unauthorized users to interact with sensitive functions or data.
Front-Running: This exploits the transparency of pending transactions. Attackers observe a pending transaction, craft a transaction with a higher gas price, and get it included in the next block first, gaining an unfair advantage.
Oracle Manipulation: Smart contracts often rely on external data sources (oracles). Attackers can compromise oracle integrity, manipulating data fed to the contract. Using multiple independent oracles and decentralized oracle networks can mitigate this risk.
These vulnerabilities underscore the importance of rigorous security analysis and testing during smart contract development and deployment.
@@ 2.2 Smart Contract Security Auditing
Static Analysis Tools: Tools like Mythril [31] and Slither [32] analyze contract source code to identify potential vulnerabilities. They perform symbolic execution and taint analysis to detect patterns associated with common vulnerabilities.
Dynamic Analysis Tools: Tools like Manticore [33]and Echidna [34] execute contracts with various inputs to uncover runtime errors. They use fuzzing and symbolic execution techniques to explore different execution paths and identify potential issues.
Formal Verification: This approach uses mathematical techniques to rigorously prove the correctness of a contract’s code against a formal specification. Tools like KEVM [35] and CertiK’s DeepSEA have been developed for formal verification of smart contracts [36].
@@
2.3 LLMs for Code Repair
LLMs, trained on vast datasets of code, have shown impressive capabilities in code repair tasks [37]. They can learn to understand and generate code that adheres to specific programming paradigms and best practices. However, applying LLMs to smart contract code repair presents unique challenges due to the specific syntax, semantics, and security considerations of languages like Solidity and Move.
@@ 3 Data Collection and Analysis Methodology
This research employs a multi-faceted approach to investigate the security of smart contracts, focusing on both Solidity and Move programming languages. The methodology encompasses the collection and analysis of three distinct datasets: Solidity-based, Move-based source code. Each dataset serves a specific purpose in addressing the research questions and contributing to a comprehensive understanding of smart contract vulnerabilities.
@@ 3.2.1 Solidity-Based Dataset
This dataset comprises a collection of vulnerable Solidity smart contracts sourced from the ”Not-So-Smart Contracts” repository curated by Trail of Bits [38]. This repository is renowned for its comprehensive set of contracts that intentionally exhibit a variety of common vulnerabilities. These vulnerabilities were chosen for inclusion because of their prevalence in real-world decentralized applications and their representation of typical errors during smart contract development. The dataset contains 60 vulnerable contracts, encompassing 8 distinct vulnerability categories. These categories’ distribution is shown in Table 1.
@@
Reentrancy $1515$ $25.025.0$
Integer Overflow/Underflow $1010$ $16.716.7$
Denial of Service (DoS) $88$ $13.313.3$
Access Control Issues $1212$ $20.020.0$
Uninitialized Storage Pointers $55$ $8.38.3$
Tx.origin Misuse $33$ $5.05.0$
Timestamp Dependency $44$ $6.76.7$
Gas Limit and Out-of-Gas Vulnerabilities $33$ $5.05.0$
@@ 3.2.2 Move-Based Dataset (Source Code)
This dataset encompasses the source code of 92 real-world Move projects, comprising 652 individual modules. These projects were part of Aptos [20], Sui [21], and Starcoin [19]. These projects span various application domains, as depicted in Table 2. The total number of Move projects is 92, and the total number of Move modules within these projects is 652.
@@
Decentralized Finance $4141$ $44.644.6$
Token $2222$ $23.923.9$
Bridge $1818$ $19.619.6$
Library $33$ $3.33.3$
Infrastructure $33$ $3.33.3$
Other $55$ $5.45.4$
@@
For both the Move-based datasets, we utilize Song et al’s [14] work to compare the vulnerability detection part. While the prior work is directed towards detection, the same dataset helps us compare Smartify’s performance on both detection and repair.
@@ 3.3 Evaluation of Smartify
Smartify is designed with two core functionalities: detecting and repairing unsafe coding patterns in smart contracts. To rigorously evaluate these capabilities, we utilize the previously described datasets, encompassing both Solidity and Move code. The evaluation process focuses on the complete output of Smartify rather than individual components, reflecting its nature as an integrated solution for smart contract security. Performance is measured using the Pass@1 score.
@@ 3.4 Agent-Based Code Repair Process for Smart Contracts
Our approach leverages a multi-agent system inspired by established software development methodologies but specifically tailored for the automated repair of Solidity and Move smart contracts. This system employs five specialized agents: an Auditor, an Architect, a Code Generator, a Refiner, and a Validator. The process incorporates a self-refinement loop and a final validation step, ensuring a high degree of accuracy and security. Each agent plays a distinct role in a structured workflow, detailed below.
3.4.1 Agent Roles and Responsibilities
  • Auditor: This agent is the cornerstone of the security analysis. It is fine-tuned on a comprehensive corpus of Solidity and Move code documentation, encompassing syntax, semantics, and best practices. Furthermore, it is safety-aligned using a classifier adapted from Google’s Responsible AI toolkit. This classifier has been meticulously modified to enforce language-specific rules and safe coding practices, effectively preventing the generation of unsafe or unsupported code constructs.
    This alignment is of paramount importance. For Move, it ensures that generated code strictly adheres to the conventions of the target blockchain (e.g., Sui or Aptos). It prevents the accidental introduction of elements from one Move variant into another or the inclusion of unsupported Rust paradigms. This is crucial because Move, while derived from Rust, has its own unique features and limitations. For Solidity, it enforces established security best practices and prevents the generation of code patterns known to be vulnerable.
    The Auditor’s primary responsibility is to meticulously scan the input smart contract code (either Solidity or Move) to identify potential vulnerabilities and unsafe patterns. It’s secondary, yet vital, role is to serve as the final validator of the repaired code.
  • Architect: This agent receives the output from the Auditor, which includes a detailed report of identified vulnerabilities and unsafe code segments. The Architect’s role is to devise a high-level strategic plan for addressing these issues. This plan does not involve generating code directly. Instead, it outlines the necessary modifications, refactoring, and improvements required to rectify the identified problems. This plan serves as a comprehensive blueprint for the Code Generator, guiding the code repair process.
  • Code Generator: This agent is a general-purpose code LLM. Its strength lies in its ability to leverage Retrieval-Augmented Generation (RAG) from two distinct data stores, one dedicated to Solidity and the other to Move. These data stores contain a collection of best practices and relevant documentation for respective programming language.
@@ 3.4.1 Agent Roles and Responsibilities
  • Refiner: This agent’s role is to enhance the quality of the code produced by the Code Generator. It achieves this through a process of iterative self-refinement, essentially acting as its own critic. The Refiner uses the same underlying LLM as the Code Generator but with a different prompt that focuses on improving the code quality based on best practices and potential improvements that it might detect from a higher level.
  • Validator: This agent acts as a final checkpoint in the process. It re-employs the Auditor agent to re-evaluate the code after the refinement stage. The Validator’s objective is to ensure that all previously identified vulnerabilities have been adequately addressed and that no new vulnerabilities have been introduced during the repair and refinement process.
Figure 1: Architecture of Smartify.

### v2

# Securing Smart Contract Languages with a Unified Agentic Framework for Vulnerability Repair in Solidity and Move

@@ Abstract
The rapid growth of the blockchain ecosystem and the increasing value locked in smart contracts necessitate robust security measures. While languages like Solidity and Move aim to improve smart contract security, vulnerabilities persist. This paper presents Smartify, a novel multi-agent framework leveraging Large Language Models (LLMs) to automatically detect and repair vulnerabilities in Solidity and Move smart contracts. Unlike traditional methods that rely solely on vast pre-training datasets, Smartify employs a team of specialized agents working on different specially fine-tuned LLMs to analyze code based on the underlying programming concepts and language-specific security principles. We evaluated Smartify on a dataset for Solidity and a curated dataset for Move, demonstrating its effectiveness in fixing a wide range of vulnerabilities. Our experimental results show that Smartify (Gemma2+Codegemma) achieves state-of-the-art performance, surpassing existing LLMs and even enhancing the capabilities of general-purpose models, such as Llama 3.1. Notably, Smartify can incorporate language-specific knowledge, such as the nuances of Move, without requiring massive language-specific pretraining datasets. This work offers a detailed analysis of the performance of various LLMs on smart contract repair, highlighting the strengths of our multi-agent approach and providing a blueprint for developing more secure and reliable decentralized applications in the growing blockchain landscape. We also provide a detailed description to extend the proposed technology to other similar use cases.
@@ 1 Introduction
Smart contracts, self-executing agreements with terms directly written into code, have emerged as a cornerstone of blockchain technology nath2014web; ray2023web3. Their ability to automate transactions and eliminate intermediaries has led to widespread adoption in various sectors, including finance, supply chain management, and healthcare zheng2018blockchain; karanjai2021conditional; kaleem2021event. However, the increasing complexity of smart contracts has given rise to a growing concern: security vulnerabilities vacca2021systematic. These vulnerabilities, often stemming from coding errors or design flaws, can be exploited by malicious actors, leading to significant financial losses and damage to the reputation of blockchain projects.
@@
The financial implications of smart contract vulnerabilities are substantial. Reports indicate that cumulative losses from attacks against Ethereum smart contracts alone have exceeded USD 3.1 billion by 2023 li2023smart. In the DeFi space, an estimated \$9.04 billion has been stolen due to vulnerabilities wronka2023financial. Notable incidents like the DAO hack of 2016, resulting in a \$55 million loss popper2016hacking, and the Poly Network hack in 2021, where over \$600 million was stolen polyhack, underscore the critical need for robust security measures.
@@
Traditional security auditing methods, although essential, often face limitations in terms of accuracy and scalability. This has spurred the exploration of automated techniques for vulnerability detection 10.1145/3238147.3238177; wang2020contractwardand repair, with Large Language Models (LLMs) emerging as a promising solution joshi2023repair. LLMs trained on vast datasets of code can learn to understand and generate code that adheres to specific programming paradigms and best practices. However, most of the tools available for smart contract safety are language-specific, especially Solidity dannen2017solidity. For other languages, existing tools often require scanning of compiled bytecode song2024empirical.
@@
Apart from Solidity, Move blackshear2019move has gained significant traction due to its strong focus on security. Its cutting-edge features, including a custom data type for secure operations, robust access controls via Move modules, and unique memory safety features blackshear2022move, have been particularly noteworthy. Moreover, the Move Prover, a native security framework, provides an additional layer of protection dill2022fast. Notably, several prominent blockchain platforms, such as Starcoin starcoin, Aptos devaptos, and Sui blackshear2024sui, have already adopted Move.
@@
However, despite its promising architecture, the real-world security performance of Move modules remains largely untested. Unlike Solidity-based smart contracts, which have been extensively studied through empirical research and surveys, there is a scarcity of research focused specifically on Move modules. Although some methodologies have been proposed for identifying defects in Move modules or conducting formal verification keilty2022model; park2024securing, and empirical analysis song2024empirical, a significant knowledge gap persists. Specifically, large-scale investigations into the frequency of defects in real-world Move modules and identifying and repairing potential vulnerabilities are lacking, highlighting the need for further research in this area.
This paper introduces Smartify, a framework that moves beyond naive LLM application by proposing a novel paradigm for structured, automated repair. We introduce Smartify not merely as a "multi-agent framework" but as a structured, collaborative system that mimics the workflow of an expert human security audit team. This contrasts with existing LLM-powered tools like ContractTinker, which utilize a linear Chain-of-Thought (CoT) reasoning process to break down the repair task. Smartify’s novelty lies in its process-centric design, where specialized agents (Auditor, Architect, Code Generator, Refiner, Validator) assume distinct roles within a delegative and iterative workflow. The Architect agent, for instance, does not simply reason about the next step; it formulates a comprehensive repair strategy that is then executed by other agents. Crucially, the Refiner-Validator duo establishes a feedback loop for iterative quality assurance, a feature essential for generating trustworthy patches.
@@
In this paper, our aim is to answer the following research questions related to software engineering using AI agents and the landscape of complex smart contract reasoning.
@@
  • RQ4: Can LLMs repair the bugs and fix the vulnerability?
  • RQ5: Does the proposed post-training framework be generalizable to a larger set of pre-trained LLMs?
The main contributions of our work are the following:
  • We introduce a novel role-based multi-agent architecture for smart contract repair that models the collaborative workflow of a security audit team, enhancing structured reasoning over monolithic LLM approaches.
  • We propose a rigorous methodology combining specialized fine-tuned models for deep vulnerability analysis, Retrieval-Augmented Generation (RAG) for language-specific context, and an iterative self-refinement loop for ensuring patch quality.
  • We conduct the first multi-faceted empirical evaluation of an LLM-based repair tool for both Solidity and Move, using metrics that assess not only correctness but also exploit mitigation effectiveness, semantic preservation, and code quality.
  • We present a comprehensive ablation study that systematically dissects our framework to quantify the distinct contribution of each architectural component to the overall repair performance.
@@ 2 Related Work
This section provides a critical review of the existing landscape in smart contract security, positioning Smartify relative to traditional analysis techniques and emerging LLM-based repair methodologies.
2.1 Traditional Smart Contract Security Analysis
Automated security analysis for smart contracts has traditionally been dominated by static, dynamic, and formal verification techniques. Each approach offers distinct advantages but also possesses inherent limitations, particularly when confronted with complex, logic-based vulnerabilities.
@@ 2.2 Smart Contract Security Auditing
Static Analysis Tools: Tools like Mythril muellerfile and Slither feist2019slither analyze contract source code to identify potential vulnerabilities. They perform symbolic execution and taint analysis to detect patterns associated with common vulnerabilities.
Dynamic Analysis Tools: Tools like Manticore mossberg2019manticore and Echidna grieco2020echidna execute contracts with various inputs to uncover runtime errors. They use fuzzing and symbolic execution techniques to explore different execution paths and identify potential issues.
Formal Verification: This approach uses mathematical techniques to rigorously prove the correctness of a contract’s code against a formal specification. Tools like KEVM hildenbrandt2018kevm and CertiK’s DeepSEA have been developed for formal verification of smart contracts zhong2020move.
@@
2.3 LLM-Powered Smart Contract Repair: A Taxonomy
The application of LLMs to automated program repair is a rapidly advancing field, with several distinct methodologies emerging.
2.3.1 Monolithic LLM Approaches
Early efforts in this domain involved using large, general-purpose LLMs like GPT-3 with intricate, zero-shot or few-shot prompts to generate patches. While demonstrating feasibility, these approaches often lack the domain-specific knowledge required for the high-stakes environment of smart contracts. They are prone to generating syntactically correct but semantically flawed or insecure code, as they lack a deep, ingrained understanding of blockchain-specific security paradigms blackshear2022move; bobadilla2025automated.
2.3.2 Chain-of-Thought (CoT) and Static Analysis Integration
To address the limitations of monolithic models, more structured approaches have been developed. A prominent example is ContractTinker, a tool designed for real-world smart contract repair wang2024contracttinker. ContractTinker employs a Chain-of-Thought (CoT) mechanism to guide an LLM through a sequence of reasoning steps: vulnerability localization, analysis, and patch generation. To ground the LLM’s reasoning and mitigate hallucination, it integrates static analysis techniques, including dependency analysis and program slicing, to provide relevant context from audit reports and the source code itself.
While ContractTinker represents a significant advancement by imposing a logical structure on the repair process, its CoT approach remains a fundamentally linear and sequential reasoning pipeline. In contrast, Smartify’s architecture is delegative and iterative. The Architect agent formulates a comprehensive, high-level plan, which is then delegated to specialized agents for execution. Furthermore, the explicit Refiner-Validator loop introduces a crucial mechanism for feedback and iterative quality improvement, a feature not explicitly detailed in the ContractTinker workflow. This architectural distinction moves beyond a simple chain of thought to a collaborative problem-solving process.
2.3.3 Multi-Agent Systems for Code Tasks
The concept of using multiple LLM-based agents to collaborate on complex tasks has gained traction in the broader software engineering domain, with applications in areas like code translation and generation huang2024opencoder; jiang2018contractfuzzer. These systems leverage the principle of specialization, assigning different roles or sub-tasks to individual agents to achieve a more robust and accurate outcome than a single agent could.
While Smartify aligns with this general trend, its novelty lies in its domain-specific agent roles tailored explicitly for the vulnerability repair workflow. Instead of a generic "translator" or "coder" agent, Smartify’s agents embody the distinct functions of a human security audit team: the Auditor for analysis, the Architect for strategic planning, the Code Generator for implementation, and the Refiner/Validator for quality assurance. This specialization allows for a more nuanced and effective approach to the highly specific and critical task of securing smart contracts.
@@ 3 Data Collection and Analysis Methodology
This research employs a multi-faceted approach to investigate the security of smart contracts, focusing on both Solidity and Move programming languages. The methodology encompasses collecting and analyzing two distinct datasets: Solidity-based, Move-based source code. Each dataset serves a specific purpose in addressing the research questions and contributing to a comprehensive understanding of smart contract vulnerabilities.
@@ 3.2.1 Solidity-based Dataset
This dataset comprises a collection of vulnerable Solidity smart contracts sourced from the "Not-So-Smart Contracts" repository curated by Trail of Bits githubGitHubCryticnotsosmartcontracts. This repository is renowned for its comprehensive set of contracts that intentionally exhibit a variety of common vulnerabilities. These vulnerabilities were chosen for inclusion because of their prevalence in real-world decentralized applications and their representation of typical errors during smart contract development. The dataset contains 60 vulnerable contracts, encompassing 8 distinct vulnerability categories. Table 1 shows these categories’ distribution.
@@
Reentrancy $15$ $25.0$
Integer Overflow/Underflow $10$ $16.7$
Denial of Service (DoS) 88 $13.3$
Access Control Issues $12$ $20.0$
Uninitialized Storage Pointers 55 $8.3$
Tx.origin Misuse 33 $5.0$
Timestamp Dependency 44 $6.7$
Gas Limit and Out-of-Gas Vulnerabilities 33 $5.0$
@@ 3.2.2 Move-Based Dataset (Source Code)
This dataset encompasses the source code of 92 real-world Move projects, comprising 652 individual modules. These projects were part of Aptos devaptos, Sui blackshear2024sui, and Starcoin starcoin. These projects span various application domains, as depicted in Table 2. The total number of Move projects is 92, and the total number of Move modules within these projects is 652.
@@
Decentralized Finance $41$ $44.6$
Token $22$ $23.9$
Bridge $18$ $19.6$
Library 33 $3.3$
Infrastructure 33 $3.3$
Other 55 $5.4$
@@
For both the Move-based datasets, we utilize Song et al.’s song2024empirical work to compare the vulnerability detection part. While the prior work is directed towards detection, the same dataset helps us compare Smartify’s performance on both detection and repair.
@@ 3.3 Evaluation of Smartify
Smartify is designed with two core functionalities: detecting and repairing unsafe coding patterns in smart contracts. We utilize the previously described datasets to evaluate these capabilities rigorously, encompassing both Solidity and Move code. The evaluation process focuses on the complete output of Smartify rather than individual components, reflecting its nature as an integrated solution for smart contract security. Performance is measured using the Pass@1 score.
@@ 3.4 Agent-Based Code Repair Process for Smart Contracts
Our approach leverages a multi-agent system inspired by established software development methodologies but tailored explicitly for the automated repair of Solidity and Move smart contracts. This system employs five specialized agents: an Auditor, an Architect, a Code Generator, a Refiner, and a Validator. The process incorporates a self-refinement loop and a final validation step, ensuring high accuracy and security. Each agent plays a distinct role in a structured workflow, detailed below.
Figure 1: Agentic architecture of Smartify.
  • Auditor: This agent is the cornerstone of security analysis. It is fine-tuned on a comprehensive corpus of Solidity and Move code documentation, encompassing syntax, semantics, and best practices. Furthermore, it is safety-aligned using a classifier adapted from Google’s Responsible AI toolkit. This classifier has been meticulously modified to enforce language-specific rules and safe coding practices, effectively preventing the generation of unsafe or unsupported code constructs.
    This alignment is of paramount importance. For Move, it ensures that the generated code strictly adheres to the conventions of the target blockchain (e.g., Sui or Aptos). It prevents the accidental introduction of elements from one Move variant into another or the inclusion of unsupported Rust paradigms. This is crucial because Move, derived from Rust, has unique features and limitations. For Solidity, it enforces established security best practices and prevents the generation of code patterns known to be vulnerable.
    The Auditor’s primary responsibility is to meticulously scan the input smart contract code (either Solidity or Move) to identify potential vulnerabilities and unsafe patterns. Its secondary, yet vital, role is to serve as the final Validator of the repaired code.
  • Architect: This agent receives the output from the Auditor, which includes a detailed report of identified vulnerabilities and unsafe code segments. The Architect’s role is to devise a high-level strategic plan for addressing these issues. This plan does not involve generating code directly. Instead, it outlines the necessary modifications, refactoring, and improvements to rectify the identified problems. This plan is a comprehensive blueprint for the Code Generator, guiding the code repair process.
  • Code Generator: This agent is a general-purpose code LLM. Its strength lies in its ability to leverage Retrieval-Augmented Generation (RAG) from two distinct data stores, one dedicated to Solidity and the other to Move. These data stores contain a collection of best practices and relevant documentation for the respective programming languages.
@@
  • Refiner: This agent’s role is to enhance the code quality produced by the Code Generator. It achieves this through iterative self-refinement, essentially acting as its critic. The Refiner uses the same underlying LLM as the Code Generator but with a different prompt that focuses on improving the code quality based on best practices and potential improvements that it might detect from a higher level.
  • Validator: This agent acts as a final checkpoint. After the refinement stage, it re-employs the Auditor agent to re-evaluate the code. The Validator’s objective is to ensure that all the previously identified vulnerabilities have been adequately addressed and that no new vulnerabilities have been introduced during the repair and refinement process.